OpenStudy (jchick):

Someone please help!! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (kamibug):

What are your options? :)

OpenStudy (jchick):

a. y + 3 = -6(x + 2) b. y + 2 = 6(x + 3) c. y - 3 = -6(x - 2) d. y - 2 = 6(x - 3)

OpenStudy (kamibug):

These are in point-slope form. \[y - y_1 = m(x - x_1)\] where (x_1, y_1) point and m is the slope. The point you are given is (-2, -3) and the m is -6. Can you plug those into the equation? :)

OpenStudy (jchick):

I don't know how to

OpenStudy (jchick):

Can you help me do that?

OpenStudy (jchick):

Ok

OpenStudy (jchick):

That is not one of my options though

OpenStudy (kamibug):

x_1 is -2 and y_1 is -3. and m is -6. There are minus signs in the equation and the points are negative so when you plug them in it will make the + sign. :)

OpenStudy (kamibug):

So which do you think it is?

OpenStudy (jchick):

Ok

OpenStudy (jchick):

It is either a or b

OpenStudy (jchick):

I don't know which one though

OpenStudy (jchick):

It is b right

OpenStudy (kamibug):

A is correct. :D Notice the slope in b is 6 instead of -6, so that one's wrong. :P Good job!

OpenStudy (jchick):

Ok thanks!

OpenStudy (jchick):

Can you help me with one more?

OpenStudy (kamibug):

Sure:)

OpenStudy (jchick):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

OpenStudy (jchick):

y = -5x - 1

OpenStudy (jchick):

Right?

OpenStudy (kamibug):

That's perfect! ^_^

OpenStudy (jchick):

Ok thanks

OpenStudy (kamibug):

Np!